OpenStudy (anonymous):

solve z^2+2jz+14-8j+0, where j is imaginary

OpenStudy (anonymous):

That should be = 0**

OpenStudy (kainui):

Is z real or complex?

OpenStudy (kainui):

Assuming z is complex, you can either make the substitution that \[\LARGE z=a+bi\] or \[\LARGE z=re^{i \theta}\] and try solving it this way. I chose the first way. This gives you an equation with a bunch of real and imaginary stuff added together. You can separate the two equations though into the real and imaginary parts since just like a basket full of apples and oranges, you can take all the oranges out and the equality in the basket of apples remains unchanged.

OpenStudy (perl):

is the plus sign = ?

OpenStudy (perl):

z^2+2jz+14-8j=0

OpenStudy (perl):

why dont you use quadratic formula ?

OpenStudy (perl):

fyi, you can edit your question

OpenStudy (perl):

Using quadratic formula I get z = 1 + 3j z = -1 -5j

OpenStudy (anonymous):

Yeah i rearranged it and used the quadratic formula, thanks gang!!!